OpenStudy (anonymous):

Write an equation for the line that passes through (3,1) and (6,4).

OpenStudy (anonymous):

Find the slope using the slope formula: \[\frac{Y _{2} - Y _{1}}{X _{2} - X _{1}}\] Next use the point-slope formula: \[y - y _{1} = m (x - x _{1})\]and plug in either set of coordinations for the y and x values.

OpenStudy (anonymous):

4 - 1 = 3 6 - 3 = 3 m = 1 y - 1 = 1 (x - 3) y = x - 2